# Implicit Personalization in Language Models: A Systematic Study

### v1

@@ 2.4.2 Sign-test for nominal responses
    Compute the sign of the $S_{i}-\tau$.
@@
    Count the signs, $n_{+}$ is the number of $+1$ signs and $n_{-}$ is the number of $-1$ signs
@@
    Under $\mathcal{H}_{0}$, the test statistic $T$ follows a binomial distribution $T\sim\text{Bin}(n_{-}+n_{+},0.5)$.
@@
    The p-value is calculated as $7P(T\leq n+-)=\sum_{k=0}^{n_{-}}\binom{n}{k}(0.5)^{k}(0.5)^{n-k}$
2.4.3 Multiple Hypothesis Testing
If the background $B$ is a binary variable that takes value from $\{b_{0},b_{1}\}$, we just need to apply once the above testing method. However, if there are more background values $\{b_{0},b_{1},\dots,b_{|\mathcal{B}|}\}$, we need to run the test for each pair $(b_{i},b_{j})$ from $\{b_{0},b_{1},\dots,b_{|\mathcal{B}|}\}$, which leads to the multiple testing problem. To control for the Type I errors (i.e., false positives to identify IP), we adjust the significant level $\alpha$ by the Bonferroni procedure Bonferroni (1936), which divides it by the number of tests, here $C_{2}^{|\mathcal{B}|}$. Each test $\mathcal{H}_{0}^{(ij)}$, is rejected if
$p^{(ij)}\leq\alpha/C_{2}^{|\mathcal{B}|}.$ (7)
@@ 3.2 Principles to Reason about the Ethicality of IP (for Human Designers)
We suggest future work to discuss IP on a case-by-case basis, and set up a community-wide guideline. To prepare for such advancement, we regard this paper as a pioneer study, where we will introduce several case studies to show the complexity in the ethical implications of IP.
@@ 4 Overview of Three Case Studies
As highlighted before, although the mathematical formulation in Section 2 describes the syntax of the IP problem (so that we can answer “does IP exist?”), the subsequent moral reasoning of IP (to answer “is IP good?”) requires the semantics of it, namely what exact value the application scenario $a$ and type of the background $b$ take. In this section, we introduce three meticulously designed case studies with the goal of introducing the diversity behind this problem.
@@ 4.0.0.0.2 Our Three Cases
  • Case Study 2: Education Disparity (e.g., do LLMs vary their answer’s quality when knowing the user identity is African American?)
@@ 4.0.0.0.3 Structure of Each Case Study
Given the above motivations, we systemize the procedure for each case study as follows. (Step 1) For each application scenario $a$ and the corresponding background $b$, we begin by addressing the nature of the problem and its impact. (Step 2) Next, we identify a very simple operationalization of a valid sub-instance of it, by introducing (i) proxies of $b$, (ii) the space of text inputs $\bm{\mathcal{X}}_{i}$ corresponding to a certain user background $b_{i}$, (iii) smart techniques to generate the style-transferred text inputs $\bm{\mathcal{X}}_{j}$ for each other user background $b_{j}$, and (iv) designing the distance metric $\Delta$ suited for the application $a$. (Step 3) Finally, we report the test results to answer whether IP exists in this case.
@@ 5.1.0.0.2 Background $\mathcal{B}$ and Its Proxies
To design a valid sub-instance of culture-specific QA, we contrast the American English speaker user background as $b_{0}$, with the British English speaker user background as $b_{1}$. As mentioned in our desideratum (D4), we aim at a simple operationalization when designing the test cases, which these two variants enable, as they have a well-studied set of vocabulary differences. Also mentioned in the design spirit, our work does not aim at experimental completeness to include all possible cultural variants, but the theoretical rigor.
@@ 5.2.0.0.1 Collecting the Questions for $\bm{\mathcal{X}}_{i}$
We collect questions with distinct answers depending on whether the user aligns with the American English-speaking or British English-speaking culture. Namely, given a generic question $\bm{q}$, there is an American response $\bm{y}_{0}^{*}$, and a British response $\bm{y}_{1}^{*}$. To this end, we introduce our AmbrQA [Uncaptioned image] dataset, which supplies the subjective questions from GlobalOpinionQA Durmus et al. (2023) with the same number of objective, fact-based questions that we collect. See Table 2 for some example questions in our dataset, and see Section D.1 for data collection details.
@@ 5.2.0.0.2 Simple Style Transfer across $\bm{\mathcal{X}}_{0}$ and $\bm{\mathcal{X}}_{1}$
5.2.0.0.3 Adapting the Distance Metric $d$
To apply our hypothesis testing method, we design a distance function $d:\mathcal{Y}\times\mathcal{Y}\rightarrow[0,100\%]$ to score the differences of each pair of responses, across all answer types. Briefly, for multiple-choice questions, we record the classification accuracy; for scale values, we report the absolute difference; and for free-text answers, we use GPT-4 to score their similarity following Deshpande et al. (2023). See details in Section D.3.
@@ 6.1.0.0.1 Application $a$
For our second application, we look at a scenario where IP is undesired, such as education disparity. To make the setup well-defined and easy to evaluate, we consider the educational essay generation task, where the task input is an essay prompt (see examples in Table 5), and the output is essay writing for which we can evaluate the quality.
@@ 6.2.0.0.1 Collecting the Original Data $\bm{\mathcal{X}}_{0}$
We collect a dataset of 518 essay prompts in the SAE style as our $\bm{\mathcal{X}}_{0}$ data. To ensure the officiality of the dataset, we look into standard English tests such as GRE and TOEFL, compiling all the 338 available GRE writing prompts by the Educational Testing Service (ETS), [^2] https://ets.org/and collecting 180 TOEFL essay prompts from a list of educational websites. See data collection details in Section E.1.
@@ 6.2.0.0.2 Text Style Transfer to Get $\bm{\mathcal{X}}_{1}$ and $\bm{\mathcal{X}}_{2}$
Finally, for this essay generation task, we query LLMs with the prompt “Please write an essay of approximately 300-350 words in response to the following question: \$\backslash n$Question: [prompt]\$\backslash n$Begin your essay below:”.
@@ 6.3 Findings
We show the average essay score difference $\mu_{\Delta}$ and $p$-value for each pair of styles in Table 7. Overall, IP is present only in Alpaca, indicating that Alpaca has a significant difference in the essay qualities it generate across the user’s English styles. In contrast other models, such as GPT-4, the Llama2 series, and the Vicuna series, all do not demonstrate disparity in the quality. It implies that those models do not pay more attention to the correlation of certain English styles in the input and the quality of the generated text.
@@ 7.1.0.0.1 Application $a$
Lastly, we look into an application where IP has complicated moral implications. We investigate the case of echo chamber Garimella et al. (2018), [^3] I.e., amplifying misinformation given users’ previous susceptibility.which implies a trade-off between user satisfaction at the moment, and long-term effect on the user and society. This task checks whether LLMs provide a correct or wrong fact given indications of users’ past exposure to misinformation.
@@ 7.1.0.0.2 Background $\mathcal{B}$ and Its Proxies
Centric to this task is the users prior exposure to misinformation, for which we use $b_{0}$ to represent users who believe in a previous correct fact, and $b_{1}$ for users who believe in a previous wrong fact.
@@ 7.2 Operationalization
7.2.0.0.1 Collecting the Correct and Wrong Facts to Compose $\bm{\mathcal{X}}_{0}$ and $\bm{\mathcal{X}}_{1}$
@@ 7.3 Findings
The results in Table 9 are concerning – most LLMs demonstrate significant echo chamber behavior. GPT-4, Llama2 series, and Vicuna series all decreases their accuracy by over 7 points when seeing the user’s prior belief in a wrong fact. Adding the GPT-3.5-Instruct model to supply more observations, we find that the models that are less influenced by users prior belief, e.g., Alpaca and GPT-3.5-Instruct, do not necessarily have more “integrity,” but actually perform poorly in the baseline setting at the first place, with only 20+accuracy.
@@ Acknowledgment
This material is based in part upon work supported by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039B; by the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645; by a National Science Foundation award (#2306372); by a Swiss National Science Foundation award (#201009) and a Responsible AI grant by the Haslerstiftung. The usage of OpenAI credits are largely supported by the Tübingen AI Center. Zhijing Jin is supported by PhD fellowships from the Future of Life Institute and Open Philanthropy, as well as the travel support from ELISE (GA no 951847) for the ELLIS program.
@@ D.3 Score Function
D.3.0.0.1 Adapting the Distance Metric $\Delta$
To apply our hypothesis testing method, we design a distance function $\Delta:\mathcal{Y}\times\mathcal{Y}\rightarrow[0,100\%]$ to score the differences of each pair of responses, across all answer types. Briefly, for multiple-choice questions, we record the classification accuracy; for scale values, we report the absolute scalar difference; and for free-text answers, we use an LLM to score their similarity following the latest practice Deshpande et al. (2023), and rescale the results to $[0,1]$. See details below.
@@ G.0.0.0.1 Inferring User Demographics
Previous literature has demonstrated the presence of implicit personal traits in human-written data (McPherson et al., 2001; Holmes and Meyerhoff, 2008; Eisenstein et al., 2014; Flek, 2020; Chen et al., 2023). Experiments have been conducted using NLP models to infer personal traits such as gender (Burger et al., 2011; Fink et al., 2012; Ciot et al., 2013; Sap et al., 2014), age (Rao et al., 2010; Nguyen et al., 2011; Morgan-Lopez et al., 2017), ethnicity (Preoţiuc-Pietro and Ungar, 2018; Abid et al., 2021), geolocation (Han et al., 2012; Graham et al., 2014), and personality (Wei et al., 2017; Gjurković and Šnajder, 2018; Mehta et al., 2020). However, many of these studies lack a clear mathematical formulation for demographic detection and focus only on limited demographic groups and data sources (Wang et al., 2019; Murray and Durrell, 1999). We propose a systematic approach that employs hypothesis testing to infer IP in LLMs.
@@ G.0.0.0.2 Responsible Use of Implicit Personalization
IP in LLMs presents both opportunities and challenges (Flek, 2020; Raharjana et al., 2021). Inferred IP can enhance NLP tasks by tailoring LLM’s responses (Hovy, 2015; Benton et al., 2016; Sasaki et al., 2018; Zeng et al., 2019). However, IP also introduces potential risks. For instance, the presence of IP can lead to implicit gender, religion, and racial biases (Bolukbasi et al., 2016; Garg et al., 2018; Wang et al., 2020; Cheng et al., 2021; Arora et al., 2023; Das et al., 2023; He et al., 2024). Even the choice of language can influence the exhibited cultural values Arora et al. (2023); Das et al. (2023). Additionally, issues such as sycophancy may arise, where models disproportionately flatter users (Sharma et al., 2023; Wei et al., 2023), and fail to keep their stance when confronted with incorrect arguments (Wang et al., 2023). Through three case studies, our work illustrates both the benefits (Case 1) and risks (Case 2 and 3) of IP, paving the way for future research to explore and address these complexities.

### v2

@@ 2.4.2 Sign-Test for Non-Interval Responses
    Count the number of similarity scores smaller than $M_{0}$ (denoted as $n_{-}$), and the number of non-zero similarity scores (denoted as $n$)
@@
    The test statistic $x$ equals $n_{-}$.
@@
    Under $\mathcal{H}_{0}$, the test statistic $X$ follows a binomial distribution $X\sim\text{Bin}(n,0.5)$, denote its cumulative density function as $cdf(\cdot,n,0.5)$.
@@
    The p-value is calculated as $P(X\geq n_{-})=1-cdf(n_{-},n,0.5)$.
If the derived p-value is smaller than the significance level $\alpha$, then $H_{0}$ is rejected, the median of the similarity score is significantly smaller than the pre-defined threshold $M_{0}$, suggesting the existence of IP.
2.5 Multiple Hypothesis Testing
If the background $B$ is a binary variable that takes value from $\{b_{0},b_{1}\}$, we just need to apply once the above testing method. However, if there are more background values $\{b_{0},b_{1},\dots,b_{|\mathcal{B}|}\}$, we need to run the test for each pair $(b_{i},b_{j})$ from $\{b_{0},b_{1},\dots,b_{|\mathcal{B}|}\}$, which leads to the multiple testing problem. To control for the Type I errors (i.e., false positives to identify IP), we adjust the significant level $\alpha$ by the Bonferroni procedure Bonferroni (1936), which divides it by the number of tests, here (${|\mathcal{B}|\choose 2}$. Each test $\mathcal{H}_{0}^{(ij)}$, is rejected if
$p^{(ij)}\leq\frac{\alpha}{{|\mathcal{B}|\choose 2}}~{}.$ (9)
@@ 3.2 Principles to Reason about the Ethicality of IP (for Human Designers)
We suggest future work to discuss IP on a case-by-case basis, and set up a community-wide guideline. We treat this paper as an introductory study, introducing several case studies to show the complexity in the ethical implications of IP.
@@ 4 Overview of Three Case Studies
Although the mathematical formulation in Section 2 describes the syntax of the IP problem (so that we can answer “does IP exist?”), the subsequent moral reasoning of IP (to answer “is IP good?”) requires the semantics of it, namely what exact value the application scenario $a$ and type of the background $b$ take. In this section, we introduce three meticulously designed case studies with the goal of introducing the diversity behind this problem.
@@ 4.0.0.0.2 Our Three Cases
  • Case Study 2: Education Disparity (e.g., Do LLMs exhibit bias by providing lower-quality answers to questions in non-standard varieties of English?)
@@ 4.0.0.0.3 Structure of Each Case Study
Given the above motivations, we systematize the procedure for each case study as follows. (Step 1) For each application scenario $a$ and the corresponding background $b$, we begin by addressing the nature of the problem and its impact. (Step 2) Next, we identify a very simple operationalization of a valid sub-instance of it, by introducing (i) proxies of $b$, (ii) the space of text inputs $\bm{\mathcal{X}}_{i}$ corresponding to a certain user background $b_{i}$, (iii) smart techniques to generate the style-transferred text inputs $\bm{\mathcal{X}}_{j}$ for each other user background $b_{j}$, and (iv) designing the distance metric $\Delta$ suited for the application $a$. (Step 3) Finally, we report the test results to answer whether IP exists in this case.
@@ 5.1 Motivation and Problem Setup
5.1.0.0.1 Motivation
Recent research has increasingly focused on the influence of cultural and linguistic variations on language models (Chen et al., 2024). Findings indicate that these variations can significantly impact response accuracy and user satisfaction (Luo et al., 2019; Ma et al., 2021; Cho et al., 2022; Huang et al., 2023; Li et al., 2024). Building on this line of work, we apply our framework to examine the model’s cultural adaptability through IP in this case study.
@@ 5.1.0.0.3 Background $\mathcal{B}$ and Its Proxies
To design a valid sub-instance of culture-specific QA, we contrast the American English speaker’s background as $b_{0}$, with the British English speaker’s background as $b_{1}$. As mentioned in our desideratum (D4), we aim at a simple operationalization when designing the test cases, which these two variants enable, as they have a well-studied set of vocabulary differences. Also mentioned in the design spirit, our work does not aim at experimental completeness to include all possible cultural variants, but the theoretical rigor.
@@ 5.2.0.0.1 Collecting the Questions for $\bm{\mathcal{X}}_{i}$
We collect questions with distinct answers depending on whether the user aligns with the American English-speaking or British English-speaking culture. Namely, given a generic question $\bm{q}$, there is an American response $\bm{y}_{0}^{*}$, and a British response $\bm{y}_{1}^{*}$. To this end, we introduce a dataset AmbrQA [Uncaptioned image]. AmbrQA [Uncaptioned image] consists of a mix of subjective questions from GlobalOpinionQA Durmus et al. (2023) and an equal number of objective, fact-based questions that we collect. See Table 2 for some example questions in our dataset, and see Section D.1 for data collection details.
@@ 5.2.0.0.2 Simple Style Transfer across $\bm{\mathcal{X}}_{0}$ and $\bm{\mathcal{X}}_{1}$
5.2.0.0.3 Adapting the Similarity Metric $s$
To apply our hypothesis testing method, we design a similarity function $s:\mathcal{Y}\times\mathcal{Y}\rightarrow[0,100\%]$ to score the similarities of each pair of responses, across all answer types. The similarity scores are calculated as follows: for multiple-choice questions, we record the classification accuracy; for scale values, we report the absolute similarity; and for free-text answers, we use GPT-4 to score their similarity following Deshpande et al. (2023). See details in Section D.3.
@@ 6.1 Motivation and Problem Setup
6.1.0.0.1 Motivation
Studies have revealed that LLMs exhibit biases as educators when explicitly informed about a student’s demographic background, such as race or social status (Kotek et al., 2023; Weissburg et al., 2024). However, there remains limited research on whether these biases persist when explicit demographic details are absent, and the user’s background must be inferred from their utterances (Warr et al., 2024; Kantharuban et al., 2024; Eloundou et al., 2024; Gonen et al., 2024).
@@ 6.1.0.0.2 Application $a$
Building on this, our second application explores a context in which such biases are particularly problematic with IP, namely education disparity. To make the setup well-defined and easy to evaluate, we consider the educational essay generation task, where the task input is an essay prompt (see examples in Table 5), and the output is essay writing for which we can evaluate the quality.
@@ 6.2.0.0.1 Collecting the Original Data $\bm{\mathcal{X}}_{0}$
We collect a dataset of 518 essay prompts in the SAE style as our $\bm{\mathcal{X}}_{0}$ data. We look into standard English tests such as GRE and TOEFL, compiling all the 338 available GRE writing prompts by the Educational Testing Service (ETS), [^2] https://ets.org/and collecting 180 TOEFL essay prompts from a list of educational websites. See data collection details in Section E.1.
@@ 6.2.0.0.2 Text Style Transfer to Get $\bm{\mathcal{X}}_{1}$ and $\bm{\mathcal{X}}_{2}$
Finally, for this essay generation task, we query LLMs with the prompt “Your task is to write an essay (about 300-350 words) in response to the following prompt.\$\backslash n$ Essay Prompt: [prompt]”.
@@ 6.3 Findings
We show the average essay score difference $\mu_{\Delta}$ and $p$-value for each pair of styles in Table 7. Overall, IP is present only in Alpaca, indicating that Alpaca has a significant difference in the essay qualities it generate across the user’s English styles. In contrast other models, Llama2-7B,-13B, Vicuna-7B, -13B, and GPT-4 do not demonstrate disparity in the response quality. This implies that these models do not provide disparaged pedagogical help for students who speak a non-standard variety of English.
@@ 7.1 Motivation and Problem Setup
7.1.0.0.1 Motivation
Sycophancy, where models tend to agree with users to increase satisfaction, has been widely recognized as a major shortcoming of current LLMs (Sharma et al., 2023b; Wei et al., 2023; Sun et al., 2024; Park et al., 2024). A related but distinct phenomenon is the echo chamber effect, which not only reinforces users’ beliefs but also tends to amplify and perpetuate pre-existing views that may involve misinformation or biased perspectives (Garimella et al., 2018). Despite its significant societal implications, the echo chamber effect remains relatively underexplored.
@@ 7.1.0.0.2 Application $a$
To delve deeper, we examine an application where this phenomenon could lead to complicated moral implications. We investigate whether an LLM can create an echo chamber for its user Garimella et al. (2018), [^3] I.e., amplifying misinformation given users’ previous susceptibility.which implies a trade-off between user satisfaction at the moment, and long-term effect on the user and society. This task checks whether LLMs provide a correct or wrong fact given indications of users’ past exposure to misinformation.
@@ 7.1.0.0.3 Background $\mathcal{B}$ and Its Proxies
Central to this task is the users prior exposure to misinformation, for which we use $b_{0}$ to represent users who believe in a previous correct fact, and $b_{1}$ for users who believe in a previous wrong fact.
@@ 7.2 Operationalization
7.2.0.0.1 Collecting the Correct and Incorrect Facts to Compose $\bm{\mathcal{X}}_{0}$ and $\bm{\mathcal{X}}_{1}$
@@ 7.3 Findings
The results in Table 9 are unsettling – most LLMs act as an echo chamber for their users by providing them with potential misinformation. Llama2-7B,-13B, Vicuna-7B,-13B, and GPT-4 all decrease their accuracy by over 7 points when seeing the user’s prior belief in a wrong fact. Adding the GPT-3.5-Instruct model to supply more observations, we find that the models that are less influenced by users prior belief, e.g., Alpaca and GPT-3.5-Instruct, are not more resilient to implicit personalization but perform poorly in the baseline setting at the first place, with only 20+accuracy.
@@ Acknowledgment
This material is based in part upon work supported by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039B; by the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645; by a National Science Foundation award (#2306372); by a Swiss National Science Foundation award (#201009) and a Responsible AI grant by the Haslerstiftung. The usage of OpenAI credits are largely supported by the Tübingen AI Center.
@@ D.3 Score Function
D.3.0.0.1 Adapting the Similarity Metric $s$
To apply our hypothesis testing method, we design a similarity function $s:\mathcal{Y}\times\mathcal{Y}\rightarrow[0,100\%]$ to score the similarities of each pair of responses, across all answer types. Briefly, for multiple-choice questions, we record the classification accuracy; for scale values, we report the absolute scalar similarity; and for free-text answers, we use an LLM to score their similarity following the latest practice Deshpande et al. (2023), and rescale the results to $[0,1]$. See details below.
@@ G.0.0.0.1 Inferring User Demographics
Previous literature has demonstrated the presence of implicit personal traits in human-written data (McPherson et al., 2001; Holmes and Meyerhoff, 2008; Eisenstein et al., 2014; Flek, 2020; Chen et al., 2023). Experiments have been conducted using NLP models to infer personal traits such as gender (Burger et al., 2011; Fink et al., 2012; Ciot et al., 2013; Sap et al., 2014), age (Rao et al., 2010; Nguyen et al., 2011; Morgan-Lopez et al., 2017), ethnicity (Preoţiuc-Pietro and Ungar, 2018; Abid et al., 2021), geolocation (Han et al., 2012; Graham et al., 2014), and personality (Wei et al., 2017; Gjurković and Šnajder, 2018; Mehta et al., 2020). However, many of these studies lack a clear mathematical formulation for demographic detection and focus only on limited demographic groups and data sources (Wang et al., 2019; Murray and Durrell, 1999). We propose a systematic approach that employs hypothesis testing to infer IP in LLMs. Another recent research by Gonen et al. (2024) on semantic leakage reveals that language models may inadvertently incorporate irrelevant information from the prompt into generation patterns, which highlights another layer of bias and unintended personalization behaviors.
@@ G.0.0.0.2 Responsible Use of Implicit Personalization
IP in LLMs presents both opportunities and challenges (Flek, 2020; Raharjana et al., 2021). Inferred IP can enhance NLP tasks by tailoring LLM’s responses (Hovy, 2015; Benton et al., 2016; Sasaki et al., 2018; Zeng et al., 2019). However, IP also introduces potential risks. For instance, the presence of IP can lead to implicit gender, religion, and racial biases (Bolukbasi et al., 2016; Garg et al., 2018; Wang et al., 2020; Cheng et al., 2021; Arora et al., 2023; Das et al., 2023; He et al., 2024; Kantharuban et al., 2024; Eloundou et al., 2024). Even the choice of language can influence the exhibited cultural values Arora et al. (2023); Das et al. (2023). Additionally, issues such as sycophancy may arise, where models disproportionately flatter users (Sharma et al., 2023a; Wei et al., 2023), and fail to keep their stance when confronted with incorrect arguments (Wang et al., 2023). Through three case studies, our work illustrates both the benefits (Case 1) and risks (Case 2 and 3) of IP, paving the way for future research to explore and address these complexities.